Correction for problem 1. it should read as:
Solve using addition to eliminate one variable.
1. -y=2x, 3x+y=20
Thanks again... You rock Alan!!

1. -y=2x, 3x+y=20
2. 4x+2y=6, 2x+y=3
Solve using Substitution to eliminate one variable.
3. y+5x=17, 4x-3y=6
4. y=-x+3, 2x=-2y+10
thanks for any help offered.
Also corrected problem 4 to its proper form... So still looking for help with Problems 1 and 4. Thank you for the help!!

Found 2 solutions by Alan3354, sweetness4g:
Answer by Alan3354(30993)

solve using addition to eliminate one variable.
1. -y=2x, 3x+y+20
The 2nd eqn has no equal sign.
------------------
2. 4x+2y=6, 2x+y=3
The 1st eqn is 2 times the 2nd.
--> dependent, an infinite # of solutions.
---------------------
Solve using Substitution to eliminate one variable.
3. y+5x=17, 4x-3y=6
y = -5x+17
Sub for y in the 2nd eqn
4x - 3(-5x+17) = 6
19x - 51 = 6
x = 3
-------
y = -5x+17
y = 2
-------------
4. y+-x+3, 2x=-2y+10
1st eqn has no equal sign.


Answer by sweetness4g(1)